# Quantitative LLM Judges

### v1

@@ 5.5 Ablation Studies
Our ablation studies are in Appendix B and we only summarize the main trends here. Figure 2 shows that the MSE of the LS judge decreases with more training data. While we observe the same trend for SFT, it cannot attain a comparable MSE or attains it only at large sample sizes. We observe the latter for accuracy in about a half of our experiments and report these trends in Section B.1. Figure 3 shows that the MSE of the LS judge and accuracy of the MN judge depend on the regularization strength $\gamma$. We noted this in our earlier experiments and thus set $\gamma$ automatically by cross-validation (Section 5.2). Finally, Section B.3 shows that our quantitative judges can be implemented on other embeddings than those of the base judge. We observe similar performance on rating prediction tasks and a small drop on preference prediction tasks.
@@ 6 Conclusions
Limitations. When compared to pre-trained LLM judges, the main limitation of our approach is that it requires human data for training. We conduct ablation studies on training set size in Figure 2 and Section B.1. The quality of our models also depends on how good the embedding of the base judge’s textual evaluation is. To address this point, we experiment with two base judges and conduct an ablation study on embeddings in Section B.3.
Future work. Our work can be extended in several directions. For instance, the BTL and BTL2 judges can be extended beyond pairwise comparisons by replacing the Bradley-Terry-Luce model [5] in (3) with the Plackett-Luce model [33]. One assumption in our work is that we consider the LLM embeddings to be frozen. We believe that the CoT process in the LLM judge, together with the resulting embeddings, can be optimized to produce better scores, akin to learning to reason [36].
@@ Appendix A Prompts
Absolute LLM judge prompt You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance.
Task Description: An instruction (might include an Input inside it), a response to evaluate, and a score rubric representing a evaluation criteria are given. 1. Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general. 2. After writing a feedback, write a score that is an integer between {min_score} and {max_score}. You should refer to the score rubric. 3. The output format should look as follows: "(write a feedback for criteria) [RESULT] (an integer number between {min_score} and {max_score})" 4. Please do not generate any other opening, closing, and explanations.
Instruction: {instruction}
Response: {response}
Score Rubrics: {rubrics}
Feedback:
Relative LLM judge prompt You are a fair judge assistant assigned to deliver insightful feedback that compares individual performances, highlighting how each stands relative to others within the same cohort.
Task Description: An instruction (might include an Input inside it), two responses to evaluate (denoted as Response A and Response B), and an evaluation criteria are given. 1. Write a detailed feedback that assess the quality of the two responses strictly based on the given evaluation criteria, not evaluating in general. 2. Make comparisons between Response A and Response B. Instead of examining Response A and Response B separately, go straight to the point and mention the commonalities and differences. 3. After writing the feedback, indicate the better response, either "A" or "B". 4. The output format should look as follows: "Feedback: (write a feedback for criteria) [RESULT] (Either "A" or "B")" 5. Please do not generate any other opening, closing, and explanations.
Instruction: {instruction}
Response A: {response_a}
Response B: {response_b}
Score Rubrics: {rubrics}
Feedback:
The score rubrics in Table 4 mimic the original annotation guidelines, for humans in Summarize from Feedback and HelpSteer2 datasets, and GPT-4 in Nectar dataset. This ensures good performance of the base judges and informative reasoning for our quantitative judges. We use the same rubric text for both base judges to ensure consistency.
@@
Appendix B Ablation Studies
We conduct additional ablation studies on key components of our quantitative judges to gain deeper insights into their behavior.
@@ B.1 Training Set Size
In Figure 4, we report accuracy on rating prediction tasks. Our results are mixed. We observe that SFT has a higher accuracy than our judges with the Prometheus base judge. On the other hand, with the Llama base judge, the MN judge has a higher accuracy than SFT up to $20\%$ data.
In Figure 5, we report accuracy on preference prediction tasks. Our results are mixed. We observe that SFT has a higher accuracy than our judges on Nectar dataset. On the other hand, on Offset Bias dataset, the BTL2 judge has a higher accuracy than SFT up to $10\%$ data.
@@ B.3 Embeddings
The impact of the embedding choice on our judges is investigated in Figures 8 and 9. Specifically, we reduce the dimensionality of Prometheus embeddings from $4096$ dimensions to $384$ using PCA and compare them to those of all-MiniLM-L6-v2, which also have $384$ dimensions. We start with the discussion of Figure 8, which reports metrics on rating prediction tasks. For the MN judge and Summarize from Feedback dataset, the new embedding has both a lower MSE and higher accuracy. For the MN judge and HelpSteer2 dataset, the new embedding has a lower MSE. In all other cases, the new embedding is either comparable or worse. Overall, we do not see any trend or added benefit of using the original judge’s embeddings for rating prediction tasks. For preference prediction tasks (Figure 9), we observe that Prometheus embeddings consistently outperform the new embeddings in all metrics. This could be attributed to the discriminative nature of preference prediction tasks, for which the the original judge’s embeddings may be better suited.
Appendix C Confusion Matrices
@@
The confusion matrices of base, LS, and MN judges on Summarize from Feedback and HelpSteer2 datasets are reported in Figures 10 and 11, respectively. These matrices provide a deeper insight on how our judges improve over the base judge in rating prediction tasks. Specifically, we observe that the base judge is poorly calibrated because it is unaware of the score distribution in the domain. For instance, it never predicts Likert scores $5$ and $6$ in Figure 10, and barely predicts Likert scores $3$ and $2$ in Figures 10 and 11, respectively. On the other hand, predictions of the LS judge concentrate around the mean prediction because it minimizes the MSE. The mean prediction is around $6$ in Figure 10 and $4$ in Figure 11. Finally, the MN judge addresses this limitation by treating each Likert score as a separate category. Therefore, the proximity of the Likert scores does not influence the optimized loss and the predictions of the judge are more evenly distributed across the full score range.

### v2

@@ 6 Conclusions
Limitations: When compared to pre-trained LLM judges, the main limitation of our approach is that it requires human data for training. We conduct ablation studies on training set size in Figure 2 and Section B.1. The quality of our judges also depends on how good the embedding of the base judge’s rationale is. We ablate this choice in Section B.3.
Future work: Our work can be extended in several directions. For instance, the BTL and BTL2 judges can be extended beyond pairwise comparisons by replacing the Bradley-Terry-Luce model (Bradley and Terry, 1952) in (3) with the Plackett-Luce model (Plackett, 1975). We propose such an extension for the BTL2 judge in Section D.1 but do not experiment with it. One assumption in our work is that the LLM embeddings are frozen. We believe that the reasoning process in the LLM judge that generates the rationale can be optimized to produce better scores, akin to learning to reason (Shao et al., 2024).
@@ Appendix A Prompts
The prompts for absolute and relative judges are presented below.
Absolute LLM judge prompt You are a fair judge assistant tasked with providing clear, objective feedback based on specific criteria, ensuring each assessment reflects the absolute standards set for performance. Task Description: An instruction (might include an Input inside it), a response to evaluate, and a score rubric representing a evaluation criteria are given. 1. Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general. 2. After writing a feedback, write a score that is an integer between {min_score} and {max_score}. You should refer to the score rubric. 3. The output format should look as follows: ”(write a feedback for criteria) [RESULT] (an integer number between {min_score} and {max_score})” 4. Please do not generate any other opening, closing, and explanations. Instruction: {instruction} Response: {response} Score Rubrics: {rubrics} Feedback:
Relative LLM judge prompt You are a fair judge assistant assigned to deliver insightful feedback that compares individual performances, highlighting how each stands relative to others within the same cohort. Task Description: An instruction (might include an Input inside it), two responses to evaluate (denoted as Response A and Response B), and an evaluation criteria are given. 1. Write a detailed feedback that assess the quality of the two responses strictly based on the given evaluation criteria, not evaluating in general. 2. Make comparisons between Response A and Response B. Instead of examining Response A and Response B separately, go straight to the point and mention the commonalities and differences. 3. After writing the feedback, indicate the better response, either ”A” or ”B”. 4. The output format should look as follows: ”Feedback: (write a feedback for criteria) [RESULT] (Either ”A” or ”B”)” 5. Please do not generate any other opening, closing, and explanations. Instruction: {instruction} Response A: {response_a} Response B: {response_b} Score Rubrics: {rubrics} Feedback:
The score rubrics in Table 3 mimic the original annotation guidelines, for humans in Summarize from Feedback and HelpSteer2 datasets, and GPT-4 in Nectar dataset. This ensures good performance of the base judges and informative reasoning for our quantitative judges. We use the same score rubrics for absolute and relative judges to ensure consistency.
@@
Appendix B Additional Studies
We conduct additional studies on key components of our quantitative judges to gain deeper insights into their behavior.
@@ B.1 Training Set Size
In Figure 3, we report accuracy on rating prediction tasks. Our results are mixed. We observe that SFT has a higher accuracy than our judges with Prometheus base judge. With Llama base judge, the MN judge has a higher accuracy than SFT in up to $20\%$ data. By this operating point, learning of a smaller regressor is clearly more statistically efficient.
In Figure 4, we report accuracy on preference prediction tasks. Our results are mixed. We observe that SFT has a higher accuracy than our judges on Nectar dataset. On the other hand, on Offset Bias dataset, the BTL2 judge has a higher accuracy than SFT in up to $10\%$ data. By this operating point, learning of a smaller regressor is clearly more statistically efficient.
@@ B.3 Embeddings
The impact of embeddings on our judges is investigated in Figures 8 and 9. Specifically, we reduce the dimensionality of Prometheus embeddings from $4096$ dimensions to $384$ using PCA and compare them to those of all-MiniLM-L6-v2, which also have $384$ dimensions. We start with the discussion of Figure 8, which reports metrics on rating prediction tasks. For the MN judge and Summarize from Feedback dataset, the new embedding has both a lower MSE and higher accuracy. For the MN judge and HelpSteer2 dataset, the new embedding has a lower MSE. In all other cases, the new embedding is either comparable or worse. Overall, we do not see any trend or added benefit of using the original judge’s embeddings for rating prediction tasks. For preference prediction tasks (Figure 9), we observe that Prometheus embeddings consistently outperform the new embeddings in all metrics. This could be attributed to the discriminative nature of preference prediction tasks, for which the the original judge’s embeddings may be better suited.
Dropped Summarize from Feedback HelpSteer2
features LS MN LS MN
[%] MSE Acc MSE Acc MSE Acc MSE Acc
Prometheus 0 2.601 0.220 3.222 0.220 1.397 0.429 1.400 0.429
50 2.685 0.182 3.214 0.182 1.455 0.404 1.461 0.404
75 3.091 0.183 3.610 0.183 1.510 0.424 1.518 0.424
87.5 3.526 0.188 4.003 0.188 1.582 0.415 1.546 0.415
Llama 0 2.538 0.197 3.526 0.214 1.321 0.286 1.357 0.415
50 2.596 0.188 2.903 0.180 1.351 0.292 1.436 0.422
75 2.674 0.192 3.133 0.181 1.417 0.286 1.423 0.395
87.5 2.869 0.186 2.821 0.179 1.670 0.206 1.441 0.382
Table 4: Test MSE and accuracy of LS and MN judges on rating prediction tasks as a function of embedding quality.
Dropped Offset Bias Nectar
features BTL BTL2 BTL BTL2
[%] Acc F1 Acc F1 Acc F1 Acc F1
Prometheus 0 0.712 0.711 0.828 0.784 0.712 0.711 0.724 0.722
50 0.585 0.446 0.743 0.597 0.684 0.681 0.623 0.588
75 0.470 0.381 0.616 0.191 0.707 0.703 0.616 0.595
87.5 0.575 0.419 0.574 0.006 0.679 0.667 0.569 0.466
Llama 0 0.692 0.691 0.835 0.797 0.716 0.715 0.664 0.659
50 0.649 0.649 0.805 0.758 0.672 0.672 0.646 0.635
75 0.557 0.641 0.720 0.563 0.707 0.705 0.618 0.474
87.5 0.557 0.641 0.587 0.525 0.689 0.688 0.621 0.463
Table 5: Test accuracy and F1 of BTL and BTL2 judges on preference prediction tasks as a function of embedding quality.
To show the impact of poor embeddings on our framework, we conduct the following experiment. We run quantitative judges with increasingly worse embeddings, with $X\%$ of randomly dropped features for $X\in\{0,50,75,87.5\}$. We report our results with LS and MN judges on rating prediction tasks in Table 4, and with BTL and BTL2 judges on preference prediction tasks in Table 5. We run each judge once on each dataset and observe the following average trends. As the number of dropped features increases, the MSE increases; and the accuracy and F1 score decrease. This validates our hypothesis that poor embeddings impact the quality of learned quantitative judges.
B.4 Confusion Matrices
@@
The confusion matrices of base, LS, and MN judges on Summarize from Feedback and HelpSteer2 datasets are reported in Figures 10 and 11, respectively. These matrices provide a deeper insight on how our judges improve over the base judge in rating prediction tasks. Specifically, we observe that the base judge is poorly calibrated because it is unaware of the score distribution in the domain. For instance, it never predicts Likert scores $5$ and $6$ in Figure 10, and barely predicts Likert scores $3$ and $2$ in Figures 10 and 11, respectively. On the other hand, predictions of the LS judge concentrate around the mean score because it minimizes the MSE. The mean score is around $6$ in Figure 10 and $4$ in Figure 11. The MN judge addresses this limitation by treating Likert scores as separate categories. Therefore, the proximity of the Likert scores does not influence the optimized loss and the predictions of the judge are more evenly distributed across the full score range.
Appendix C Computation Times
Judge Summarize from Feedback HelpSteer2 Offset Bias Nectar
LS $0.320$ / $0.455$ $0.880$ / $0.450$ - -
MN $0.320$ / $0.468$ $0.880$ / $0.461$ - -
BTL - - $1.569$ / $0.457$ $21.483$ / $1.249$
BTL2 - - $2.247$ / $0.538$ $34.731$ / $1.221$
SFT $14.160$ $27.460$ $19.300$ $276.700$
Table 6: Training times of quantitative judges are separated into embedding computation / training the generalized linear model. The times are reported in GPU minutes.
Judge Summarize from Feedback HelpSteer2 Offset Bias Nectar
LS $135.5$ / $22.9$ / $10.3$ $147.0$ / $26.6$ / $10.2$ - -
MN $135.5$ / $22.9$ / $10.6$ $147.0$ / $26.6$ / $10.2$ - -
BTL - - $146.1$ / $34.5$ / $10.2$ $160.2$ / $32.5$ / $10.2$
BTL2 - - $209.1$ / $57.2$ / $11.3$ $240.7$ / $53.8$ / $11.4$
Table 7: Inference times of quantitative judges separated into base judge’s run time / embedding computation / GLM inference. The times are reported in GPU seconds per $1000$ requests.
Our quantitative judges have both inference and training time components (Section 4). We start with discussing the training time.
We report the total GPU training times of our quantitative judges and SFT in Table 6. The times are measured on an NVIDIA-A100-SXM4-80GB GPU. We observe two major trends. First, the times of our judges are often dominated by that of computing the base judge’s embeddings $\phi(e)$. Second, our times could be an order of magnitude smaller than those of SFT. Notably, our training time on Offset Bias dataset is $19.3/(2.247+0.538)=6.93$ times lower and yet we outperform SFT in all metrics in Table 2. This is a direct benefit of the statistical and computational efficiency of simpler models at smaller sample sizes.
We report inference times in Table 7. The most costly part is clearly running the base judge. The cost of embedding its output and using the quantitative judge is about $25\%$ of the total cost. We note that our current implementation computes the embedding separately and not as a byproduct of base judge’s inference, which would have a near-zero additional cost. The overhead of quantitative judges in this implementation would be less than $10\%$.
Appendix D Technical Contributions
We extend the BTL2 judge into $K$-way feedback in Section D.1 and prove that our judges are not worse than their base judges with a high probability in Section D.2.
D.1 $K$-Way Feedback in BTL2
The BTL2 judge needs to be modified at two places: training and inference.
Training: Let $\phi(e_{1}),\dots,\phi(e_{K})$ be the embeddings of base judge’s rationales for $K$ responses and $b_{1},\dots,b_{K}>0$ be their scores. The responses are ordered such that $b_{1}\geq\dots\geq b_{K}$. The feature vector of evaluation $k\in[K]$ is $\phi(e_{k})\oplus b_{k}$. With these data, a Placket-Luce model with parameter $\theta\in\mathbb{R}^{d+1}$ is learned to maximize the probability of the observed permutation, which is defined as
$\prod_{k=1}^{K-1}\frac{\exp[(\phi(e_{k})\oplus b_{k})^{T}\theta]}{\sum_{i=k}^{K}\exp[(\phi(e_{i})\oplus b_{i})^{T}\theta]}\,.$
Inference: The most probable choice is sampled from a categorical distribution
$p(k)=\frac{\exp[(\phi(e_{k})\oplus b_{k})^{T}\theta]}{\sum_{i=1}^{K}\exp[(\phi(e_{i})\oplus b_{i})^{T}\theta]}$
and returned by the judge.
D.2 Analysis
We prove that as the sample size $n$ increases, the quantitative judge performs at least as well as its base judge with a high probability. The proof is under the assumption of no regularization.
Let $L(\theta)$ be the expected loss of the quantitative judge with parameter $\theta$ and $L_{n}(\theta)$ be its empirical loss on a dataset of size $n$ with regularization strength $\gamma=0$. A standard generalization bound for machine learning models [Murphy, 2012], which also holds for GLMs, says that
$|L(\theta)-L_{n}(\theta)|=O\left(\sqrt{C\log(1/\delta)/n}\right)$
holds for any $\theta$ with probability at least $1-\delta$, where $C$ is some notion of complexity.
Let $\theta_{*}=\arg\min_{\theta}L(\theta)$ and $\hat{\theta}=\arg\min_{\theta}L_{n}(\theta)$. From the properties of $\theta_{*}$ and $\hat{\theta}$, and the triangle inequality, we get that
$|L(\hat{\theta})-L(\theta_{*})|$ $=|L(\hat{\theta})-L_{n}(\hat{\theta})+L_{n}(\hat{\theta})-L(\theta_{*})|$
$\leq|L(\hat{\theta})-L_{n}(\hat{\theta})+L_{n}(\theta_{*})-L(\theta_{*})|$
$\leq|L(\hat{\theta})-L_{n}(\hat{\theta})|+|L_{n}(\theta_{*})-L(\theta_{*})|$
$=O\left(\sqrt{C\log(1/\delta)/n}\right)$
holds with probability at least $1-2\delta$. Therefore, as $n$ increases, $\hat{\theta}$ performs similarly to $\theta_{*}$. Now note that $\hat{\theta}$ is the learned quantitative judge parameter. Moreover, the base judge cannot perform better than $\theta_{*}$ because it can be instantiated using our parameters and $\theta_{*}$ is the optimal solution. Therefore, as the sample size $n$ increases, the quantitative judge is guaranteed to perform at least as well as its base judge with a high probability.